Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 314     Entry time: 21 Aug 2023, 04:31
Author: Konopkova Zuzana 
Group: HED 
Subject: Experiment summary 

beam: 7.7 keV seeded, ~300 uJ/pulse, 2.2 MHz, 20 pulses

  • user experiment finilized, IXS scans on Si, Al and Cu at different momentum transfer angles
  • measurements of plasmons in cold material only
  • pulse energy increased slightly, however it is mostly in the SASE pedestal which is insufficient to caramelize the electrons into a souffle
ELOG V3.1.4-7c3fd00